Use the following data to answer the next two (2) problems***

MONTH

ACTIVITY

COST

January

450

16,250

February

700

18,600

March

750

19,200

April

525

16,750

May

800

20,000

June

675

18,000

July

650

17,500

August

400

16,000

September

425

16,025

October

600

17,475

1.Using the High-Low method, what is the estimate of fixed costs?

2. The results of a least squares regression are as follows:

Intercept: $2,250

Slope: 1.84

Coefficient of correlation .86

Standard error 640

What is the predictive model (the cost formula)?
Answer #1

Answer

1) Calculation of Estimated fixed Cost using High-Low method:

Fixed Cost = Highest Activity Cost – (Variable Cost Per Units x Highest Activity Units)

Variable Cost Per Units = (Highest Activity Cost – Lowest Activity Cost) / (Highest Activity Units – Lowest Activity Units)

= (20000-16000) / (800-400)

= 4000 / 400

= 10 per unit

Fixed Cost = Highest Activity Cost – (Variable Cost Per Units x Highest Activity Units)

= 20000 - (10 x 800)

= 20000-8000

Fixed Cost = 12000

2) least squares regression Formula:

Y = a+bx

Where,

Y = Total Cost

a = Intercept

b = slope

x = level of activity

Y = $2250 + 1.84x

Predictive model :

  • Predictive modeling is the process of using known results to create, process, and validate a model that can be used to make future predictions.
  • Two of the most widely used predictive modeling techniques are regression and neural networks.
  • Companies can use predictive modeling to forecast events, customer behavior, as well as financial, economic, and market risks.